# Training Convolutional Neural Networks with the Forward-Forward algorithm

### 22 Dec 2023 v1

@@ Abstract
The recent successes in analyzing images with deep neural networks are almost exclusively achieved with Convolutional Neural Networks (CNNs). The training of these CNNs, and in fact of all deep neural network architectures, uses the backpropagation algorithm where the output of the network is compared with the desired result and the difference is then used to tune the weights of the network towards the desired outcome. In a 2022 preprint, Geoffrey Hinton suggested an alternative way of training which passes the desired results together with the images at the input of the network. This so called Forward Forward (FF) algorithm has up to now only been used in fully connected networks. In this paper, we show how the FF paradigm can be extended to CNNs. Our FF-trained CNN, featuring a novel spatially-extended labeling technique, achieves a classification accuracy of 99.0% on the MNIST hand-written digits dataset. We show how different hyperparameters affect the performance of the proposed algorithm and compare the results with CNN trained with the standard backpropagation approach. Furthermore, we use Class Activation Maps to investigate which type of features are learnt by the FF algorithm.
Keywords: Forward-Forward, keyword two, keyword three
@@ 1 Introduction
Machine learning using deep neural networks (DNN) continues to transform human live in areas as different as art (DALL-E, stable diffusion), medicine (Alpha-Fold), transport (self-driving cars, any time now), or information retrieval (ChatGPT, Bard). Here, the adjective deep refers to the number of layers of artificial neurons which can go up to the hundreds. Training these networks means to shift the weights connecting the layers from their initial, random values to values which produce the correct predictions at the output layer of the DNN. This is achieved with the help of a loss function that computes the aggregate difference between the predicted output and the accurate results, which must be known for the training examples. The algorithm behind the training is some variant of gradient descent: in each round of training each weight is shifted a bit into the direction minimizing the loss by using the derivative of the loss function with respect to that weight.
Taking the derivative of a loss function with respect to a given weight is straightforward for a single, output layer. Training the weights of the earlier layers in DNN requires the gradient computation to be performed iteratively by applying the chain rule (Rumelhart et al., 1986). This process is called backpropagation (BP). Due to its importance, the term BP is also often used loosely to refer to the entire learning algorithm including the gradient descent.
@@
Backpropagation, respectively multi-layer gradient descent, has a number of downsides: first, it requires the storage of intermediate results. Depending on the optimizer, the memory consumption of BP is up to 5 times larger than the requirement for storing the weights alone (Hugging Face Community, 2023). This becomes a problem when training large models on GPU cards with limited memory. Second, under the name neuromorphic computing there is an ongoing search of hardware alternatives to CMOS semiconductors, driven by the desire to decrease power consumption and increase processing speed (Christensen et al., 2022). On these new hardware platforms it is often impossible to implement an analog of BP, raising the need for alternative training algorithm. Finally, evolution has clearly developed learning algorithms for neural networks such as our brain. However, those algorithms seem to be quite (but maybe not completely (Lillicrap et al., 2020)) different from BP. Given the in general high performance of evolutionary solutions to problems, this raises the question if also deep learning could gain from biologically plausible alternatives to BP.
@@
Due to this limitations, there is an ongoing search for alternative training methods. The most radical approach is to abandon the loss function completely and use a new learning paradigm. Neural networks trained with variants of the locally acting Hebbian learning rule (neurons that fire together wire together) have been shown to be competitive with BP (Journé et al., 2023; Zhou, 2022). Another approach, Equilibrium Propagation (Scellier and Bengio, 2017) is a learning framework for energy based models with symmetric connections between neurons. Layer-wise Feedback Propagation (Weber et al., 2023) removes the need for a gradient computation by replacing the objective of reducing the loss with computing a reward signal from the network output and propagating that signal backward into the net.
@@
Another example of the use of local information gathered in consecutive runs to update the weights is the Forward-Forward (FF) algorithm proposed by Geoffrey Hinton in December 2022. FF training combines two ideas. First, the weights of a given layer are updated using gradients of a locally defined goodness function, which in the original paper is taken to be the sum of the squares of the activities in that layer. Second, the labels are included in the training data which allows the neurons to learn them together. In order to understand which features of the data ’vote’ for a given label, half of the dataset is comprised of labels combined with wrong images. For this negative data the weights are changed in order to minimize the goodness. In contrast, for the correctly labeled, positive data the weights are modified to maximize the goodness. Both these objectives can be achieved with a local gradient descent, with no need for BP. The term Forward Forward now refers to having two subsequent training steps, one with positive and one with negative data.
@@
Firstly, we showed how we developed our FF-trained CNN. We described the implementation of a spatial-extended labeling, crucial to preserve the label information in the image during convolution, and the mathematical steps underneathing both the FF-based learning algorithm and the inference approaches proposed by Hinton (2022). Then, we showed the results of the tests performed with our first FF-trained CNN architecture on the hand-written digits of the MNIST dataset: we obtained a classification accuracy of 99.14% for the validation set and 99.00% for the test dataset, comparable to the results of conventional backpropagation-based networks. Furthermore, we characterized the novel developed CNN by evaluating how its hyperparameters affect the network performances. Finally, we provided the results of a preliminary investigation, conducted by exploiting Class Activation Maps (CAMs), aimed to evaluate which image-based features are processed by the FF paradigm to accomplish classification task.
@@ 2.1 Spatially-extended labelling
As indicated in Figure 1 C and D, both the positive and negative datasets are obtained by harnessing this methodology. Labels are created by starting from an empty 2D Fourier space of the same size than the image. We then set a single 2D Fourier mode to a finite amplitude . The mapping between the labels and the wavelength and orientation of the corresponding mode is a matter of choice; we have tested two different options. Set 1, which is reported in Figure 2, is characterized by a combination of 4 different wave orientations (0°, 45°, 90°, 135°) and three different frequencies. Set 2, as shown in Figure 7 in the appendix, uses only one spatial frequency and obtains the different labels from ten equidistant angular orientations in the range [0°, 180°]. After the superposition of the label, the images are again normalized to the range 0 to 1. The relative contribution of the label pattern to the total intensity of the image is a hyperparameter whose influence will be described in section 3.2.3. Note that we choose our negative labels randomly, not specifically hard based on another forward pass, as it was done in Hinton (2022).
@@ 2.2 Implementation of the learning algorithm
The present study uses a network architecture composed of three consecutive FF-trained convolutional layers. All three layers contain the same number of filter matrices which is one of the hyperparameter we examined. We did not add any max pooling layers because we found that those decrease accuracy as described in section C in the appendix. All neurons use ReLU as activation function.
All results reported are obtained using the MNIST handwritten digits dataset (LeCun and Cortes (2010)). We split the original training data into 50,000 images used for training and 10,000 validation images used in the search for the optimal hyperparameters. The original 10,000 test images were only used for obtaining the final result reported in section 3.1. All reported accuracies are averaged over 5 independent runs.
Just as in the FF training of DNN, we evaluate the loss of each convolutional layer by computing the sigmoidal function of the goodness. The goodness is defined as the sum of squared layer activations , modified by subtracting a user-provided threshold . Following the code provided by Geoffrey Hinton333 The webpage of Prof. Hinton includes a demonstration code at https://www.cs.toronto.edu/ hinton/., and as confirmed by Gandhi et al. (2023), we choose to be equal to the number of neurons within that layer.
@@
We do not induce symmetry in our loss as described in Lee and Song (2023). But we took inspiration from Lorberbom et al. (2023) who found improved collaboration between layers by training them with a cumulative network loss, computed by summing the individual layer losses . Here, we exclude the loss of the first layer since it yielded better accuracy as shown in Section B in the appendix. This also aligns with Hinton’s exclusion of the first layer during the evaluation phase.
We follow the implementation of Hinton (2022) in two more aspects. First, we apply layer normalization between the individual layers. As described in Ba et al. (2016), layer normalization involves the application of the following transformation to each activation :
@@
In order to study the contribution of the individual layers we define a layer based loss and accuracy which measure only the capability to discriminate between images of the positive and negative dataset. We interpret the output of the sigmoid function as a probability, where values 0.5 indicate that the layer recognizes the image as belonging to the positive dataset. By comparing the true assignation (positive or negative) we obtain a discrimination accuracy. And by using the probability to compute a binary cross entropy, we compute a layer-specific discrimination loss.
@@ 2.3 Two ways of inference
There are two ways how an FF trained CNN or DNN can be used for inference: the linear classifier and the goodness evaluation. In the first case, the neurons of every layer except the first are fully-connected with an output layer of nodes, equal to the number of labels. The connecting weights, times in number, are trained by evaluating the neuron activations using a cross-entropy loss.
@@
where goodness is expressed as for each associated label and is again the number of all neurons, except those from the first layer.
@@ 2.4 Hardware and Software
The code for our FF trained CNN is implemented in Python using the PyTorch library. We took the code of Loewe (2023) for an FF trained, fully connected DNN as the starting point for our own efforts. Our code is available on a GitHub 444code will be available after publication in peer-reviewed journal. All analysis presented here was performed on a desktop workstation with an AMD Ryzen 9 5900X 12-Core Processor and 125 GB RAM.
@@ 3 Results
We first report the configuration which achieved the highest accuracy, then the search through the space of hyperparameters (using only our validation dataset) leading to that optimal configuration. We close by demonstrating the ability of FF trained CNN to implement Class Activation Maps, which is a method from the explainabel AI toolbox.
3.1 Performance of optimized configuration
@@
The hyperparameters optimization lead to the following configuration for the FF trained CNN: three convolution layers of each 128 filters with a kernel dimension of 7x7 pixels. Even though the goodness computation provides a better accuracy, the inference is performed for time reasons with the linear classifier approach. After training for 200 epochs with a batch size of 50, ADAM optimizer with learning rate = , and the label set 1 (Figure 2) with intensity = 35% we obtained 99.14% accuracy for the validation dataset and 99.00% for the test dataset (the latter is presently averaged only over 2 runs).
For comparison, using again a three layer CNN of constant size but trained with BP with obtained a validation accuracy of 99.13% (16 filters of 5x5 pixels, Adam optimizer with learning rate = , 200 epochs, batch size = 50).
@@
Figure 3 provides a more detailed picture of this comparison. Figure 3 A show that the performance of FF training increases monotoneously with the number of filters while BP training accuracy decreases slightly under the same conditions. The latter is most likely an effect due to increasing overfitting; its absence indicates that FF training might either be more robust against overfitting, or that it makes less efficient use of its number of trainable parameters.
@@
The confusion matrix (Figure 3 B) provides insights into the classification performance, revealing that labels 5 and 9 exhibit the least accurate classifications (lower than 98.20%), while labels 6, 7 and 1 are characterized by the highest accuracy levels (higher than 99.50%).
@@
After 200 epochs, discrimination loss of the FF trained network reaches a plateau for all considered layers (Figure 3 C). This was confirmed by a training with 750 epochs which resulted in no substantial changes in accuracy. Moreover, the network exhibited convergence, as evidenced not only by the overall accuracy value on the training data reaching close to 100% (as depicted in Figure 3 D with a green line), but also by the discrimination accuracy values achieved by each layer (red and blue lines), which both independently recognized the images of the positive and the negative datasets.
In order to investigate the effect of the goodness of the first convolutional layer, we trained the CNN which reached the best accuracy value for a second time while adding also the goodness value computed from the first layer to the total goodness. Figure 8 in the appendix reports the results, highlighting that the training of the first layer affects the fast convergence of the next layer and reduces the overall accuracy of the network by approximately .
3.2 FF-trained CNN hyperparameters characterization
The best FF-trained CNN architecture we proposed has been obtained by characterizing different parameters, which can affect the network classification performance. These analysis were performed by using a constant learning rate of , which was empirically found to be optimal for the FF-trained network in the wide range between and , over 200 epochs. Initially, we refrained from incorporating supplementary layers such as fully-connected, dropout, or pooling layers. Indeed, only in the final phase of our investigation we introduced max pooling layers to assess their impact. This choice was motivated by our intention to underscore the autonomous capability of the novel FF-trained CNN in effectively processing information gleaned from dataset images, thus aptly executing classification tasks. Moreover, no competitive results were achieved by applying max pooling layers, as more extensively described by 2
3.2.1 Filter dimensions and Batch size
Firstly, we focused on assessing critical hyperparameters that define the FF-trained CNN, including filter dimensions used during convolution, network complexity represented by the number of filters per layer, and the exploited batch size. Figure 4 B shows that more complex FF-trained architectures and 7x7 pixel filters demonstrated superior accuracy. Specifically, the 128x128x128 architecture consistently delivered the highest accuracy across all tested filter dimensions (3x3, 5x5, 7x7). Moreover, the 7x7 filter configuration yielded the best results irrespective of the number of filters employed. The 64x64x64 configuration also provided reliable results, although it slightly trailed the top-performing setup tested (128x128x128 filters of 7x7 pixel dimensions). As shown in Figure 4 A, for 5x5 and 7x7 filters better results were achieved for lower batch sizes, while the computational time needed for the training phase increased considerably. Specifically, 43.10.8 s/epoch were necessary to train the 64x64x64 network with batch sizes of 25 elements. Conversely, employing a batch size of 50 elements under the same configuration reduced the required training time to 21.90.3 s/epoch.
3.2.2 Linear Classifier vs. Goodness approach
Another aspect to investigate pertains to the inference approach. We assessed both the linear classifier and the goodness computation methods for different CNNs, as outlined in Table 1. Each method is associated with an optimal learning rate of . As shown also by Brenig and Timofte (2023), the goodness approach yielded slightly superior results, although the computational time was longer both during the training and during the validation phases. This discrepancy arises because, in the goodness method, each image in the dataset must be processed for each label in order to perform the classification task. The goodness accuracy computation proves to be valuable for datasets with a small number of labels. However, for datasets with a larger number of labels, the marginal improvements offered by this method may not be justified. Goodness can be employed in scenarios where the computational time is not a critical factor. Since preserving computing resources is a primary objective of the FF algorithm, the linear classifier is the preferred choice. Consequently, we proceeded to evaluate the best configuration by exclusively using this approach.
FF-CNN Network Linear Classifier Goodness Accuracy [%] Time/epoch [s] Accuracy [%] Time/epoch [s] 128x128x128, batch=50, filter 5x5 98.93 43.1 0.6 s 99.04 331 13 s 16x16x16, batch=50, filter 5x5 98.16 7.60.2 s 98.75 154 1 s 64x64x64, batch=50, filter 3x3 98.30 21.90.3 s 98.82 288 6 s 64x64x64, batch=200, filter 5x5 98.67 7.60.2 s 98.94 150 1 s
3.2.3 Effect of the relative labeling intensity
We explored the impact on the network performance of varying relative intensity contributions of the labeling pattern on both positive and negative datasets. We focused on architectures that exhibited the highest accuracy values: 128x128x128 and 64x64x64 filters. We tested batch sizes of 25 and 50 elements. Each architecture featured 7x7 pixel filters, a learning rate of , 200 epochs, and employed the Adam optimizer. Comprehensive results were presented in Figure 5. Optimizing label intensity led to an increase in networks accuracy by up to 0.2%. Specifically, the tested 64x64x64 architectures exhibited their highest accuracy values when the label intensity contribution was set . In contrast, the 128x128x128 networks attained their peak result with a label relative intensity of , irrespective on the elected batch size. This leads to the conclusion that the optimal value of labeling intensity, represented by , should be determined based on the specific network architecture. Conversely, when internal parameters, such as filter dimensions and batch size, are modified, there is no need to re-optimize the value of .
3.3 Class Activation Maps for explainable AI applications

### 25 Dec 2023 v2

@@ Abstract
The recent successes in analyzing images with deep neural networks are almost exclusively achieved with Convolutional Neural Networks (CNNs). The training of these CNNs, and in fact of all deep neural network architectures, uses the backpropagation algorithm where the output of the network is compared with the desired result and the difference is then used to tune the weights of the network towards the desired outcome. In a 2022 preprint, Geoffrey Hinton suggested an alternative way of training which passes the desired results together with the images at the input of the network. This so called Forward-Forward (FF) algorithm has up to now only been used in fully connected networks. In this paper, we show how the FF paradigm can be extended to CNNs. Our FF-trained CNN, featuring a novel spatially-extended labeling technique, achieves a classification accuracy of 99.16% on the MNIST hand-written digits data set. We show how different hyperparameters affect the performance of the proposed algorithm and compare the results with CNN trained with the standard backpropagation approach. Furthermore, we demonstrate that Class Activation Maps can be used to investigate which type of features are learnt by the FF algorithm.
Keywords: forward-forward, convolutional neural networks, class activation maps
@@ 1 Introduction
Machine learning using deep neural networks (DNN) continues to transform human live in areas as different as art (DALL-E, stable diffusion), medicine (Alpha-Fold), transport (self-driving cars, any time now), or information retrieval (ChatGPT, Bard). Here, the adjective deep refers to the number of layers of artificial neurons which can go up to the hundreds. Training these networks means to shift the weights connecting the layers from their initial, random values to values which produce the correct predictions at the output layer of the DNN. This is achieved with the help of a loss function that computes the aggregate difference between the predicted output and the accurate results, which must be known for the training examples. The algorithm behind the training is some variant of gradient descent: in each round of training each weight is shifted a bit into the direction minimizing the loss by using the derivative of the loss function with respect to that weight. Taking the derivative of a loss function with respect to a given weight is straightforward for a single, output layer. Training the weights of the earlier layers in DNN requires the gradient computation to be performed iteratively by applying the chain rule (Rumelhart et al., 1986). This process is called backpropagation (BP). Due to its importance, the term BP is also often used loosely to refer to the entire learning algorithm including the gradient descent.
@@
Backpropagation, respectively multi-layer gradient descent, has a number of downsides: first, it requires the storage of intermediate results. Depending on the optimizer, the memory consumption of BP is up to 5 times larger than the requirement for storing the weights alone (Hugging Face Community, 2023). This becomes a problem when training large models on GPU cards with limited memory. Second, under the name neuromorphic computing there is an ongoing search of hardware alternatives to CMOS semiconductors, driven by the desire to decrease power consumption and increase processing speed (Christensen et al., 2022). On these new hardware platforms it is often impossible to implement an analog of BP, raising the need for alternative training algorithm. Finally, evolution has clearly developed learning algorithms for neural networks such as our brain. However, those algorithms seem to be quite (but maybe not completely, see e.g. Lillicrap et al., 2020) different from BP. Given the in general high performance of evolutionary solutions to problems, this raises the question if also deep learning could gain from biologically plausible alternatives to BP.
@@
Due to these limitations, there is an ongoing search for alternative training methods. The most radical approach is to abandon the loss function completely and use a new learning paradigm. Neural networks trained with variants of the locally acting Hebbian learning rule (neurons that fire together wire together) have been shown to be competitive with BP (Journé et al., 2023; Zhou, 2022). Another approach, Equilibrium Propagation (Scellier and Bengio, 2017) is a learning framework for energy based models with symmetric connections between neurons. Layer-wise Feedback Propagation (Weber et al., 2023) removes the need for a gradient computation by replacing the objective of reducing the loss with computing a reward signal from the network output and propagating that signal backward into the net.
@@
Another example of the use of local information gathered in consecutive runs to update the weights is the Forward-Forward (FF) algorithm proposed by Geoffrey Hinton in December 2022. FF training combines two ideas. First, the weights of a given layer are updated using gradients of a locally defined goodness function, which in the original paper is taken to be the sum of the squares of the activities in that layer. Second, the labels are included in the training data which allows the neurons to learn them together. In order to understand which features of the data vote for a given label, half of the data set is comprised of labels combined with wrong images. For this negative data the weights are changed in order to minimize the goodness. In contrast, for the correctly labeled, positive data the weights are modified to maximize the goodness. Both these objectives can be achieved with a local gradient descent, with no need for BP. The term Forward Forward now refers to having two subsequent training steps, one with positive and one with negative data.
@@
This paper is structured as following: in Section 2 we will first introduce our spatial-extended labeling technique which is crucial to preserve the label information during convolution. Then we will discuss the details of our implementation of the FF-based learning and inference. The results in Section 3 start with a discussion of the optimal results we obtained on the MNIST hand written digits classification. In Section 3.2 we describe our search for the optimal hyperparameters (using validation data). Section 3.3 shows how Class Activation Maps (CAMs) can be used to get a better understanding of the features learned during training. We close in Section 4 with a short discussion.
@@ 2 Materials and Methods
This section will first discuss our new technique for labeling the positive and negative data sets, before explaining our implementation of the FF algorithm in detail.
@@ 2.1 Spatially-extended Labelling
As indicated in Figure 1 C and D, both the positive and negative data sets are obtained by harnessing this methodology. Labels are created by starting from an empty 2D Fourier space of the same size than the image. We then set a single 2D Fourier mode to a finite amplitude. The mapping between the labels and the wavelength and orientation of the corresponding mode is a matter of choice; we have tested two different options. Set 1, which is reported in Figure 2, is characterized by a combination of 4 different wave orientations (0°, 45°, 90°, 135°) and three different frequencies. Set 2, as shown in Figure 8, Appendix A, uses only one spatial frequency and obtains the different labels from ten equidistant angular orientations in the range [0°, 180°]. The results obtained by Set 2 are slightly worse, we report them in Appendix A.
After the superposition of the label, the images are normalized to the range 0 to 1. The relative contribution of the label pattern to the total intensity of the image is a hyperparameter whose influence will be described in Section 3.2.3. Note that we choose our negative labels randomly, not specifically hard based on another forward pass, as it was done in Hinton (2022).
@@ 2.2 Implementation of the Learning Algorithm
The present study uses a network architecture composed of three consecutive FF-trained convolutional layers. All three layers contain the same number of filter matrices which is one of the hyperparameter we examine. We did not add any max pooling layers because we found that those decrease accuracy as described in Appendix C. All neurons use ReLU as activation function.
All results reported are obtained using the MNIST handwritten digits data set (LeCun and Cortes (2010)). We split the original training data into 50,000 images used for training and 10,000 validation images used in the search for the optimal hyperparameters. All reported validation accuracies are averaged over 5 independent runs with standard deviations typically being 0.01%. The original 10,000 test images were only used for obtaining the final result reported in Section 3.1.
Just as in the FF training of DNN, we evaluate the loss of each convolutional layer by computing the sigmoidal function of the goodness. The goodness is defined as the sum of squared layer activations , modified by subtracting a user-provided threshold . Following the code provided by Geoffrey Hinton333Code is available at https://www.cs.toronto.edu/hinton/ffcode.zip/, and as confirmed by Gandhi et al. (2023), we choose to be equal to the number of neurons within that layer.
@@
We do not induce symmetry in our loss as described in Lee and Song (2023). But we took inspiration from Lorberbom et al. (2023) who found improved collaboration between layers by training them with a cumulative network loss, computed by summing the individual layer losses . Here, we exclude the loss of the first layer since it yielded better accuracy as shown in Appendix B, This also aligns with Hinton’s exclusion of the first layer during the evaluation phase.
We follow the implementation of Hinton (2022) in two more aspects. First, we apply layer normalization between the individual layers. As described in Ba et al. (2016), layer normalization involves the application of the following transformation to each activation as:
@@
In order to study the contribution of the individual layers we define a layer based loss and accuracy which measure only the capability to discriminate between images of the positive and negative data set. We interpret the output of the sigmoid function as a probability, where values indicate that the layer recognizes the image as belonging to the positive data set. By comparing with the true assignation (positive or negative) we obtain a discrimination accuracy. And by using the probability to compute a binary cross entropy, we compute a layer-specific discrimination loss.
@@ 2.3 Two Ways of Inference
There are two ways how an FF trained CNN or DNN can be used for inference: the linear classifier and the goodness evaluation. In the first case, the neurons of every layer except the first are fully-connected with an output layer of nodes, equal to the number of labels. The connecting weights, times in number, are trained by evaluating the neuron activations using a cross-entropy loss. If not mentioned differently, this is the default method of inference used in this paper.
@@
where the goodness is expressed as for each associated label and is again the number of all neurons, except those from the first layer.
@@ 2.4 Hardware and Software
The code for our FF trained CNN is implemented in Python using the PyTorch library (pytorch). The source code of Loewe (2023) for an FF trained, fully connected DNN was used as the starting point and our code is available on GitHub444Code will be available after publication in a peer-reviewed journal.. All analysis presented here was performed on a desktop workstation with an AMD Ryzen 9 5900X 12-Core Processor with 128 GB RAM, and an NVIDIA RTX 3080 GPU with 12 GB RAM.
@@ 3 Results
We first report the configuration which achieved the highest accuracy, then the search through the space of hyperparameters (using only our validation data set) leading to the optimal configuration. We close by demonstrating the ability of FF trained CNN to implement Class Activation Maps, which is a method from the explainable AI toolbox.
3.1 Performance of the Optimized Configuration
@@
The hyperparameters optimization leads to the following configuration for the FF trained CNN: three convolution layers of each 128 filters with a kernel dimension of 7x7 pixels. After training for 200 epochs with a batch size of 50, Adam optimizer with a learning rate = , and the label set 1 (Figure 2) with intensity = 35%, we obtain 99.20% accuracy for the validation data set and 99.16% for the test data set using the goodness approach for inference. While having a significantly shorter run time, inference with the linear classifier approach provides slightly worse results, achieving accuracy values of 99.14% and 99.00% for validation and test data sets, respectively. For comparison, using again a three layer CNN of constant size but trained with BP we obtain a validation accuracy of 99.13%. Here the search for optimal hyperparameters resulted in 16 filters of 5x5 pixels, Adam optimizer with learning rate = , 200 epochs with batch size = 50.
@@
Figure 3 provides a more detailed picture of this comparison. Figure 3 A shows that the performance of FF training increases monotonously with the number of filters per layer while BP training accuracy decreases slightly under the same conditions. The latter is most likely due to increasing overfitting; its absence indicates that FF training might either be more robust against overfitting, or that it makes less efficient use of its number of trainable parameters.
@@
The confusion matrix in Figure 3 B provides insight into the classification performance, revealing that labels 5 and 9 exhibit the least accurate classifications (lower than 98.20%), while labels 6, 7 and 1 are characterized by the highest accuracy levels (higher than 99.50%).
@@
After 200 epochs, the discrimination loss of the FF trained network reaches a plateau for all considered layers (Figure 3 C). This convergence of the training is confirmed by a training run with 750 epochs which results in no substantial changes in accuracy. Moreover, he accuracy value on the training data reaches close to 100% (depicted in Figure 3 D with a green line) for 200 epochs. The discrimination accuracy values of the layers (red and blue lines in Figure 3 D) corroborate this result. They also hint at a slightly different dynamics of the two layers 2 and 3 with layer 2 initially learning faster, but layer 3 achieving a higher faculty of discrimination on the long run.
3.2 Influence of Hyperparameters
The performance of BP trained neural networks depends on the value of the hyperparameters used during training. There is no reason to expect this to be different for FF training. This Section discusses the search we performed in order to obtain the optimal result reported in Section 3.1.
This search also included the exploration of some architectural options which are not hyperparameters in the strict sense. We limited our search to networks of three layers of convolutional filters, all with the same number of filters per layer. Moreover, we did not include max pooling layers because the tests described in Appendix C indicate that they decrease accuracy.
As optimizer we tested stochastic gradient descent and Adam with the latter consistently outperforming the former. For the learning rate of Adam we evaluated the range between and . A learning rate of yielded the best results and was therefore chosen as default. As in Section 3.1, we found that after training for 200 epochs the network had converged in all configurations tested, which made this value also our default.
3.2.1 Filter Dimensions, Layer Width, and Batch Size
First, we describe the interaction of the two architectural parameters filter dimension and number of neurons in each layer with the training parameter batch size. Figure 4 A shows that the accuracy decreases with decreasing filter size. A possible explanation is that the smaller filters have more problems identifying the wavelengths of the label waves. For 7x7 and 5x5 filters we additionally observe an increase in accuracy for smaller batch sizes. We speculate that this might be a generic feature due to the dual nature of the training data.
Figure 4 B confirms for all filter sizes the increase of accuracy with the number of filters per layer, which was first shown in Figure 3 A. Again the accuracy decreases with decreasing filter size.
3.2.2 Inference: Linear Classifier vs. Goodness Approach
As described in Section 2.2 there are too ways for making inference with a FF trained neural network. Table 1 shows that the goodness computation approach beats the linear classifier over a variety of different CNN configurations. This finding agrees with the result of Brenig and Timofte (2023).
However, the slightly superior results go together with a ten times larger computational cost of the goodness comparison because each image in the data set must be processed for each label in order to perform the classification task. Consequentially, the goodness approach might only be justified for classification task with a limited number of labels.
| FF-CNN Network | Linear Classifier | Goodness |
| Accuracy [%] | Accuracy [%] | |
| 128 filter of size 5x5, batch size 50 | 98.93 | 99.04 |
| 16 filter of size 5x5, batch size 50 | 98.16 | 98.75 |
| 64 filter of size 3x3, batch size 50 | 98.30 | 98.82 |
| 64 filter of size 5x5, batch size 200 | 98.67 | 98.94 |
3.2.3 Effect of the Relative Labeling Intensity K
In studying the effect of the relative label intensity , we focus on the hyperparameters that exhibit the highest accuracy values: 64 or 128 convolutional filters per layer, filter size 7x7, and batch sizes of 25 and 50 images. The results presented in Figure 5 show that optimizing leads to an increase in accuracy of up to 0.2%. The optimal value of does depend on layer width but not on batch size. Specifically, the 64x64x64 architectures achieves the highest accuracy for , the 128x128x128 networks for . This relatively large difference was unexpected.
3.3 Class Activation Maps